### v1

@@ 4.1 Proof in the Finite Case
Suppose that $X_{1},X_{2}$ are independent random variables, and $\|X_{i}\|_{\Psi_{2}}\leq K_{i}$, $i=1,2$. Then $\|X_{1}+X_{2}\|_{\Psi_{2}}\leq\sqrt{K_{1}^{2}+K_{2}^{2}}$.
@@
By Jensen’s inequality,
$\mathbb{E}\left[\exp(\frac{X_{1}^{2}+X_{2}^{2}}{K_{1}^{2}+K_{2}^{2}})\right]$ $=\mathbb{E}\left[\exp(\frac{X_{1}^{2}}{K_{1}^{2}+K_{2}^{2}})\right]\mathbb{E}\left[\exp(\frac{X_{2}^{2}}{K_{1}^{2}+K_{2}^{2}})\right]$ (49)
$\leq\mathbb{E}^{\frac{K_{1}^{2}}{K_{1}^{2}+K_{2}^{2}}}\left[\exp(\frac{X_{1}^{2}}{K_{1}^{2}})\right]\mathbb{E}^{\frac{K_{2}^{2}}{K_{1}^{2}+K_{2}^{2}}}\left[\exp(\frac{X_{2}^{2}}{K_{2}^{2}})\right]$ (50)
$\leq 2.$ (51)
We also need the following estimate of the covering number:
@@
For given $\sigma>0$ independent of $N$, we have
$NR_{\mu}(4\sigma^{2})+o(N)\leq\ln N(\mathcal{C},2\sqrt{N}\sigma)\leq NR_{\mu}(\sigma^{2})+o(N)$ (52)
where $o(N)$ hides constants depending on $P_{Y}$ and $\sigma$.
@@
For any $t^{N}\in\mathcal{C}$, a standard application of Sanov’s theorem (Csiszár and Körner,, 2011, Chapter 1) shows
$\ln\frac{1}{\mu^{\otimes N}(B(t^{N},\sqrt{N}\sigma))}=NR_{\mu}(\sigma^{2})+o(N)$ (53)
where the implicit constant in $o(N)$ may depend on $\mu$. Let $P(\mathcal{C},2\sqrt{N}\sigma)$ denote the packing number, i.e. the maximal number of points $\mathcal{C}$ separated by distance at least $2\sqrt{N}\sigma$. Then $N(\mathcal{C},2\sqrt{N}\sigma)\leq P(\mathcal{C},2\sqrt{N}\sigma)$ (see e.g. Van Handel, (2014)), and standard volume estimates yields $P(\mathcal{C},2\sqrt{N}\sigma)\leq\frac{1}{\min_{t^{N}\in\mathcal{C}}\mu^{\otimes N}(B(t^{N},\sqrt{N}\sigma))}$, proving the second inequality in (52). Moreover, by a volume argument we also have $N(\mathcal{C},2\sqrt{N}\sigma)\max_{t^{N}\in\mathcal{C}}\mu^{\otimes N}(B(t^{N},2\sqrt{N}\sigma))\geq\mu^{\otimes N}(\mathcal{C})=\exp(o(N))$. Then using (53) with $\sigma$ replaced by $2\sigma$, we obtain the first inequality in (52). ∎
@@
For any integer $N>0$, let $\mathcal{C}$ be the $P_{Z}$-typical set (Thomas and Joy,, 2006):
$\mathcal{C}:=\mathcal{C}_{N}:=\{z^{N}\in T^{N}\colon\|\widehat{P}_{z^{N}}-\mu\|_{1}\leq N^{-1/3}\}$ (54)
where $\widehat{P}_{z^{N}}$ denotes the empirical distribution of $z^{N}$, and $\|\widehat{P}_{z^{N}}-\mu\|_{1}:=\sum_{z\in T}|\widehat{P}_{z^{N}}(z)-\mu(z)|$ can in fact be replaced by any norms, which are equivalent up to constants due to finiteness of $T$. Moreover, let $n:=|T|$. Without loss of generality, we can identify $T$ with the standard basis vectors $e_{1},e_{2},\dots,e_{n}$ of $\mathbb{R}^{n}$, and define $Y_{i}:=X_{e_{i}}$ for each $i\in\{1,\dots,n\}$. Then the process can be represented by inner product:
$X_{z}=\left<Y,z\right>,\quad\forall z\in T.$ (55)
It is not necessary to adopt the inner product representation for this proof, but it helps making the notations clear. Suppose that $P_{YZ}\in\Pi(P_{Y},\mu)$ is arbitrary, and let $(Y^{N},Z^{N})\sim P_{YZ}^{\otimes N}$. We have
$\mathbb{E}[\left<Y,Z\right>]=\lim_{N\to\infty}\mathbb{E}[\frac{1}{N}\left<Y^{N},Z^{N}\right$>]. (56)
By Proposition 2, $(\left<Y^{N},z^{N}\right>)_{z^{N}\in\mathcal{C}}$ is a sub-Gaussian process with respect to the metric $d_{N}(z^{N},\hat{z}^{N}):=\sqrt{\sum_{i=1}^{N}d^{2}(z_{i},\hat{z}_{i})}$. Let $1_{\mathcal{C}}$ be the indicator of the event that $Z^{N}\in\mathcal{C}$. Since $\left<Y^{N},Z^{N}\right>1_{\mathcal{C}}\leq\max_{z^{N}\in\mathcal{C}}\left<Y^{N},z^{N}\right>1_{\mathcal{C}}=\max_{z^{N}\in\mathcal{C}}\left<Y^{N},z^{N}\right>-\max_{z^{N}\in\mathcal{C}}\left<Y^{N},z^{N}\right>1_{\mathcal{C}^{c}}$, we have
$\mathbb{E}[\left<Y^{N},Z^{N}\right$>] $\leq\mathbb{E}[\left<Y^{N},Z^{N}\right>1_{\mathcal{C}^{c}}]+\mathbb{E}[\max_{z^{N}\in\mathcal{C}}\left<Y^{N},z^{N}\right>]-\mathbb{E}[\max_{z^{N}\in\mathcal{C}}\left<Y^{N},z^{N}\right>1_{\mathcal{C}^{c}}]$ (57)
$\leq\mathbb{E}[\max_{z^{N}\in\mathcal{C}}\left<Y^{N},z^{N}\right>]+\mathbb{E}[\max_{z^{N},\hat{z}^{N}\in\mathcal{Z}^{N}}\left<Y^{N},z^{N}-\hat{z}^{N}\right>1_{\mathcal{C}^{c}}].$ (58)
The last term above has vanishing contribution: Indeed, note that for each $z^{n},\hat{z}^{n}\in\mathcal{Z}^{N}$, we have $\|\left<Y^{N},z^{N}-\hat{z}^{N}\right>\|_{\Psi_{2}}\leq C\sqrt{N}$ for some $C$ independent of $N$ (may depend on $P_{YZ}$). Taking the max over $\exp(O(N))$ such sub-Gaussian random variables will increase the sub-Gaussian norm by a factor of no more than $\sqrt{N}$ (Vershynin,, 2018), so
$\|\max_{z^{N},\hat{z}^{N}\in\mathcal{Z}^{N}}\left<Y^{N},z^{N}-\hat{z}^{N}\right>\|_{\Psi_{2}}\leq CN.$ (59)
Moreover, by Hoeffding’s inequality we also have $\mathbb{E}[1_{\mathcal{C}^{c}}]=\exp(-\Theta(N^{1/3}))$, so we can use Cauchy-Schwarz (or (Talagrand,, 2014, Lemma 7.8.4)) to show the second expectation in (58) is $o(N)$. Therefore, we have verified that
$\mathbb{E}[\left<Y,Z\right>]\leq\liminf_{N\to\infty}\frac{1}{N}\mathbb{E}[\max_{z^{N}\in\mathcal{C}}\left<Y^{N},z^{N}\right$>]. (60)
Furthermore, applying Dudley’s integral (see (7) and the subsequent comments) to $\mathcal{C}$ yields
$\mathbb{E}[\max_{z^{N}\in\mathcal{C}}\left<Y^{N},z^{N}\right>]\leq 24\int_{0}^{\Delta_{\rm m}}\sqrt{\ln N(\mathcal{C},2\sqrt{N}\sigma)}2\sqrt{N}d\sigma$ (61)
where $\Delta_{\rm m}:=\sup_{N\geq 1}\frac{{\rm diam}(\mathcal{C}_{N})}{2\sqrt{N}}<\infty$. Now we multiply $1/N$ to (61), take $N\to\infty$, and apply Lemma 4. We can apply dominated convergence, since $\frac{1}{N}\sqrt{\ln N(\mathcal{C},2\sqrt{N}\sigma)}2\sqrt{N}\leq\frac{2}{\sqrt{N}}\sqrt{\ln(n^{N})}=2\sqrt{\ln n}$, which is independent of $N$ and absolutely integrable on $\sigma\in[0,\Delta_{\rm m}]$. We have thus proved Theorem 1 for finite $T$. ∎
The proof of Theorem 2 will be almost identical to the proof of Theorem 1, but we will use an additional technique about the projection to the set of sequences of a given type. More precisely:
Lemma 5.
Let $\mu$ be a rational distribution on $\mathbb{R}^{n}$. Fix $N\geq 1$ such that $N\mu$ is integer. For any $y^{N}\in\mathbb{R}^{nN}$, we have
$\frac{1}{\sqrt{N}}\min_{s^{N}\colon\widehat{P}_{s^{N}}=\mu}\|y^{N}-s^{N}\|_{2}=W_{2}(\widehat{P}_{y^{N}},\mu).$ (62)
Proof.
The $\geq$ part is obvious from the definition. For the $\leq$ part, fix an arbitrary $\bar{s}^{N}\in\mathbb{R}^{nN}$ satisfying $\widehat{P}_{\bar{s}^{N}}=\mu$. Denote the supports of $\widehat{P}_{y^{N}}$ and $\mu$ be $\mathcal{A}$ and $\mathcal{B}$, respectively. Consider the following function
$F(M)$ $=\frac{1}{N}\sum_{1\leq i,j\leq N}\|y_{i}-\bar{s}_{j}\|_{2}^{2}M_{i,j}$ (63)
on $\mathcal{S}$, which is defined as the set of doubly stochastic matrices $M\in\mathbb{R}^{N\times N}$. We can see from (64) that $\inf_{M\in\mathcal{S}}F(M)\leq W_{2}^{2}(\widehat{P}_{y^{N}},\mu)$. Since $F$ is linear, by the Birkhoff–von Neumann theorem, there exists a permutation matrix $M$ achieving $\inf_{M\in\mathcal{S}}F(M)$. This permutation sends $\bar{s}^{N}$ to an $s^{N}$, achieving $\inf_{M\in\mathcal{S}}F(M)=\frac{1}{N}\|y^{N}-s^{N}\|_{2}$. This proves the $\leq$ part. ∎
@@
As in the proof of Theorem 1, we can identify $T=\{e_{1},\dots,e_{n}\}$ as the standard basis vectors, so that $X_{z}=\left<Y,z\right$>. The assumption of zero-mean process implies that $\mathbb{E}[Y]=0$. Then, shifting $Z\sim\mu$ by a deterministic vector does not change the value of $\mathbb{E}[\left<Y,Z\right$>], so we can assume without loss of generality that $\mathbb{E}[Z]=0$. Let us assume that $P_{YZ}$ is rational; the general case will then following by approximation. Define
$\mathcal{I}:=\{N\geq 1\colon\textrm{$NP_{YZ}$ is integer}\}.$ (65)
For each $N\in\mathcal{I}$, let $\mathcal{C}_{\mu}$ denote the $\mu$-type set, consisting of all sequences of type $\mu=P_{Z}$. Select
@@
sequences from $\mathcal{C}_{\mu}$ uniformly at random with replacement, and call this random set $\mathcal{A}$. Let $Y^{N}\sim P_{Y}^{\otimes N}$. Then we will show that
$\mathbb{E}[\left<Y,Z\right>]\leq\liminf_{N\to\infty}\mathbb{E}[\sup_{z^{N}\in\mathcal{A}}\frac{1}{N}\left<Y^{N},z^{N}\right$>], (67)
where the limit is for $N\in\mathcal{I}$. Indeed, set
@@
and for any $y^{N}\in\mathcal{B}$, define $\mathcal{D}(y^{N}):=\{z^{N}\colon(y^{N},z^{N})\in\mathcal{D}\}$. A standard result states that (see Lemma 3, or (Csiszár and Körner,, 2011, Chapter 1)):
$\frac{1}{N}\ln|\mathcal{D}(y^{N})|=H(Z|Y)+O(\frac{\ln N}{N}),\quad\forall y^{N}\in\mathcal{B}.$ (70)
Let $\pi$ be the projection from $\mathbb{R}^{nN}$ to $\mathcal{B}$ (i.e., map to the closest point in $\mathcal{B}$, with ties broken arbitrarily). Define $\mathcal{E}$ as the event that $\mathcal{D}(\pi(Y^{N}))\cap\mathcal{A}=\emptyset$. Then, (70) implies that
$\mathbb{P}[\mathcal{E}]=\left(1-\exp(-NI(Y;Z)+O(\ln N))\right)^{L}\leq\exp(-CN^{1/3})$ (71)
for some $C>0$ depending only on $P_{YZ}$, and for all $N\in\mathcal{I}$. From the definitions it is clear that
@@
for all $N\in\mathcal{I}$ almost surely. Since $\mathbb{E}[Z]=0$ is assumed, we see that 0 belongs to the convex hull of $\mathcal{C}_{\mu}$, therefore
$\max_{z^{N}\in\mathcal{A}}\frac{1}{N}\left<\pi(Y^{N}),z^{N}\right>\geq\max_{z^{N}\in\mathcal{A}}\frac{1}{N}\left<\pi(Y^{N}),z^{N}\right>1_{\mathcal{E}^{c}}$ (73)
almost surely. Furthermore,
@@
$\leq W_{2}(\mu,\widehat{P}_{Y^{N}})\cdot O(1)$ (76)
which converges to 0 in expectation as $N\in\mathcal{I}$ tends to infinity. Thus (67) is verified by collecting the bounds above.
@@
$\limsup_{N\to\infty}\frac{1}{N}\ln N(\mathcal{A},2\sqrt{N}\sigma)\leq R_{\mu}(\sigma^{2})\wedge I(Y;Z).$ (77)

### v3

@@ 4 Proof of the rate-distortion integral: finite case
Suppose that $X_{1},X_{2},\dots,X_{n}$ are zero-mean, independent random variables, and $\|X_{i}\|_{G}\leq K_{i}$, $i=1,2,\dots,n$. Then $\|X_{1}+X_{2}+\dots+X_{n}\|_{G}\leq\sqrt{K_{1}^{2}+K_{2}^{2}+\dots+K_{n}^{2}}$.
Let $n:=|T|$. Without loss of generality, we can identify $T$ with the standard basis vectors $e_{1},e_{2},\dots,e_{n}$ of $\mathbb{R}^{n}$, and define $Y_{i}:=X_{e_{i}}$ for each $i\in\{1,\dots,n\}$. Then the process can be represented by inner product:
$X_{z}=\left<Y,z\right>,\quad\forall z\in T.$ (37)
It is not necessary to adopt the inner product representation for the proof of Theorem 1, but it helps making the notations clear. Now let $Y^{N}\sim P_{Y}^{\otimes N}$. Recall the definitions related to types in Section 2.3. For any $\mu$ on $T$, define
$\mathcal{I}:=\{N\in\mathbb{Z}\colon\textrm{$N>0$, $N\mu$ is integer}\}$ (38)
which is nonempty if $\mu$ is rational. For $N\in\mathcal{I}$, define the type class
$\mathcal{C}_{\mu}$ $:=\{z^{N}\in T^{N}\colon\widehat{P}_{z^{N}}=\mu\}.$ (39)
The following two lemmas show that a coupling can be approximated by a matching:
Lemma 3.
Let $\mu$ be a rational distribution on $\mathbb{R}^{n}$, and let $N\in\mathcal{I}$. For any $y^{N}\in\mathbb{R}^{nN}$, we have
$\frac{1}{\sqrt{N}}\min_{s^{N}\in\mathcal{C}_{\mu}}\|y^{N}-s^{N}\|_{2}=W_{2}(\widehat{P}_{y^{N}},\mu).$ (40)
Lemma 4.
Suppose that $\mathcal{X}$ and $\mathcal{Y}$ are finite sets, $N>0$ is an integer, and $P_{XY}$ is a distribution with the property that both $NP_{X}$ and $NP_{Y}$ are integer. Then there exists $Q_{XY}$ satisfying $Q_{X}=P_{X}$, $Q_{Y}=P_{Y}$, $\max_{x,y}|Q_{XY}(x,y)-P_{XY}(x,y)|\leq\frac{1}{N}$, and that $NQ_{XY}$ is integer.
Lemma 3 has no approximation error, whereas Lemma 4 shows approximation on the level of the distribution (not just the min value) but has a $1/N$ approximation error. For our purpose, it is in fact possible to use only Lemma 4, since the $1/N$ gap vanishes in the limit. The proofs of both lemmas can be found in the appendix.
By Proposition 2, $(\left<Y^{N},z^{N}\right>)_{z^{N}\in\mathcal{C}}$ is a sub-Gaussian process with respect to the following metric on $T^{N}$:
$d_{N}(z^{N},\hat{z}^{N}):=\sqrt{\sum_{i=1}^{N}d^{2}(z_{i},\hat{z}_{i})}.$ (41)
The following observation is crucial:
Lemma 5.
For any rational $\mu$, we have
$\lim_{\mathcal{I}\ni N\to\infty}\frac{1}{N}\mathbb{E}[\max_{z^{N}\in\mathcal{C}_{\mu}}\left<Y^{N},z^{N}\right>]=\mathsf{w}(\mu),$ (42)
where $\lim_{\mathcal{I}\ni N\to\infty}$ denotes the limit along the subsequence in $\mathcal{I}$.
@@
By Lemma 3, we see that for any $N\in\mathcal{I}$,
$\frac{1}{N}\max_{z^{N}\in\mathcal{C}_{\mu}}\left<y^{N},z^{N}\right$> $=\frac{\mathbb{E}_{\widehat{P}_{y^{N}}}[\|Y\|_{2}^{2}]+\mathbb{E}_{\mu}[\|Z\|_{2}^{2}]-W_{2}^{2}(\widehat{P}_{y^{N}},\mu)}{2}.$ (43)
Define $M_{N}:=\sup_{P_{Y^{\prime}}\colon W_{2}(P_{Y},P_{Y^{\prime}})\leq N^{-\frac{1}{2n}}}\frac{\mathbb{E}_{P_{Y^{\prime}}}[\|Y\|_{2}^{2}]+\mathbb{E}_{\mu}[\|Z\|_{2}^{2}]-W_{2}^{2}(P_{Y^{\prime}},\mu)}{2}$. Since $W_{2}$ is a metric (satisfying the triangle inequality), we see that $\mathbb{E}_{P_{Y^{\prime}}}[\|Y\|_{2}^{2}]-W_{2}^{2}(P_{Y^{\prime}},\mu)$ is continuous in $P_{Y^{\prime}}$ with respect to $W_{2}$, and hence
$\lim_{\mathcal{I}\ni N\to\infty}M_{N}$ $=\sup_{P_{YZ}\in\Pi(P_{Y},\mu)}\mathbb{E}[\left<Y,Z\right$>] (44)
$=\mathsf{w}(\mu).$ (45)
Define
$\mathcal{D}$ $:=\{y^{N}\in\mathbb{R}^{nN}\colon W_{2}(P_{Y},\widehat{P}_{y^{N}})\leq N^{-\frac{1}{2n}}\}.$ (46)
By the Wasserstein law of large numbers (cf. [8]), we have $\lim_{\mathcal{I}\ni N\to\infty}\mathbb{P}[Y^{N}\in\mathcal{D}]=1$. Therefore,
$\limsup_{\mathcal{I}\ni N\to\infty}\frac{1}{N}\mathbb{E}[\max_{z^{N}\in\mathcal{C}}\left<Y^{N},z^{N}\right>1_{Y^{N}\in\mathcal{D}}]$ $=\limsup_{\mathcal{I}\ni N\to\infty}\mathbb{E}[\sup_{P_{YZ}\in\Pi(\widehat{P}_{Y^{N}},\mu)}\mathbb{E}[\left<Y,Z\right>]1_{Y^{N}\in\mathcal{D}}]$ (47)
$\leq\lim_{\mathcal{I}\ni N\to\infty}M_{N}\mathbb{E}[1_{Y^{N}\in\mathcal{D}}]=\mathsf{w}(\mu),$ (48)
where the first equality follows from Lemma 3. Similarly, if the $\sup$ in the definition of $M_{N}$ is replaced by $\inf$, then (45) still holds, but the inequality sign in (48) will be reversed, and the $\limsup$ in (47) will be replaced by $\liminf$. Hence we have
$\lim_{\mathcal{I}\ni N\to\infty}\frac{1}{N}\mathbb{E}[\max_{z^{N}\in\mathcal{C}}\left<Y^{N},z^{N}\right>1_{Y^{N}\in\mathcal{D}}]=\mathsf{w}(\mu).$ (49)
It remains to show that $\lim_{\mathcal{I}\ni N\to\infty}\frac{1}{N}\mathbb{E}[\max_{z^{N}\in\mathcal{C}_{\mu}}\left<Y^{N},z^{N}\right>1_{Y^{N}\in\mathcal{D}^{c}}]=0$, which, by the Cauchy-Schwarz inequality, would follow once we show
$\limsup_{\mathcal{I}\ni N\to\infty}\mathbb{E}\left[\left(\frac{1}{N}\max_{z^{N}\in\mathcal{C}_{\mu}}\left<Y^{N},z^{N}\right>\right)^{2}\right]<\infty.$ (50)
To show (50), first note that, due to $|T|<\infty$, we have $\|\frac{1}{N}\left<Y^{N},z^{N}\right>\|_{G}\leq\frac{C_{1}}{\sqrt{N}}$ for each $z^{N}\in\mathcal{C}$, where $C_{1},C_{2},C_{3}>0$ here are some constants independent of $N$. Since $\sqrt{\ln|\mathcal{C}_{\mu}|}\leq C_{2}\sqrt{N}$, by [43, Exercise 2.37] we have $\|\frac{1}{N}\max_{z^{N}\in\mathcal{C}_{\mu}}\left<Y^{N},z^{N}\right>\|_{G}\leq C_{3}\cdot\frac{C_{1}}{\sqrt{N}}\cdot C_{2}\sqrt{N}=C_{1}C_{2}C_{3}$. Then (50) follows since the sub-Gaussian norm dominates the square root of the second moment. ∎
We have the following estimate of the covering number $\mathsf{N}(\mathcal{C},\cdot)$ under the metric $d_{N}$:
@@
Let $\mu$ be rational, and $\mathcal{C}:=\mathcal{C}_{\mu}$. For any $\sigma>0$, we have
$\limsup_{\mathcal{I}\ni N\to\infty}\frac{1}{N}\ln\mathsf{N}(\mathcal{C},2\sqrt{N}\sigma)$ $\leq i_{\mu}(\sigma);$ (51)
$\liminf_{\mathcal{I}\ni N\to\infty}\frac{1}{N}\ln\mathsf{N}(\mathcal{C},2\sqrt{N}\sigma)$ $\geq i_{\mu}(2\sigma).$ (52)
@@
For any $t^{N}\in\mathcal{C}$, we have
$\mu^{\otimes N}(B(t^{N},\sqrt{N}\sigma))=\sum_{P_{Z\hat{Z}}\in\mathcal{S}_{N}}\mu^{\otimes N}(\mathcal{C}_{P_{Z\hat{Z}}}(t^{N}))$ (53)
where $\mathcal{S}_{N}$ denotes set of all $P_{Z\hat{Z}}\in\Pi_{\sigma}(\mu)$ such that $NP_{Z\hat{Z}}$ is integer, and $\mathcal{C}_{P_{Z\hat{Z}}}(t^{N}):=\{u^{N}\in{T}^{N}\colon\textrm{$(t^{N},u^{N})$ has type $P_{Z\hat{Z}}$}\}$. Evidently,
$\max_{P_{Z\hat{Z}}\in\mathcal{S}_{N}}\mu^{\otimes N}(\mathcal{C}_{P_{Z\hat{Z}}}(t^{N}))\leq\sum_{P_{Z\hat{Z}}\in\mathcal{S}_{N}}\mu^{\otimes N}(\mathcal{C}_{P_{Z\hat{Z}}}(t^{N}))\leq|\mathcal{S}_{N}|\max_{P_{Z\hat{Z}}\in\mathcal{S}_{N}}\mu^{\otimes N}(\mathcal{C}_{P_{Z\hat{Z}}}(t^{N}))$ (54)
where $\lim_{\mathcal{I}\ni N\to\infty}\frac{1}{N}\ln|\mathcal{S}_{N}|=0$ by Lemma 1. By (22), we have
$\left|\frac{1}{N}\ln\mu^{\otimes N}(\mathcal{C}_{P_{Z\hat{Z}}}(t^{N}))-I(Z;\hat{Z})\right|\leq c_{N}$ (55)
where $c_{N}:=\frac{2|{T}|^{2}}{N}\ln(N+1)$. Next, we show that
$\lim_{\mathcal{I}\ni N\to\infty}\max_{P_{Z\hat{Z}}\in\mathcal{S}_{N}}I(Z;\hat{Z})=i_{\mu}(\sigma).$ (56)
Indeed, the $\leq$ part is obvious. For the $\geq$ part, let $P_{Z\hat{Z}}^{*}$ be the optimal distribution in the definition of $\Pi_{\sigma}(\mu,\mu)$. By Lemma 4, there exists $P^{N}_{Z\hat{Z}}\in\mathcal{S}_{N}$ such that $\max_{z,u}|P_{Z\hat{Z}}^{*}(z,u)-P^{N}_{Z\hat{Z}}(z,u)|\leq\frac{1}{N}$. Therefore $\lim_{\mathcal{I}\ni N\to\infty}I(Z;\hat{Z})=i_{\mu}(\sigma)$ and $\lim_{\mathcal{I}\ni N\to\infty}\mathbb{E}[d^{2}(Z,\hat{Z})]=\sigma^{2}$ where $(Z,\hat{Z})\sim P^{N}_{Z\hat{Z}}$. Since $i_{\mu}(\sigma)$ is continuous with respect to $\sigma$, this establishes the $\geq$ part of (56). Now combining (54), (55) and (56), we obtain
$\left|\frac{1}{N}\ln\frac{1}{\mu^{\otimes N}(B(t^{N},\sqrt{N}\sigma))}-i_{\mu}(\sigma)\right|\leq c_{N}^{\prime}$ (57)
where $c_{N}^{\prime}>0$ is a vanishing sequence possibly depending on $\sigma$ and $\mu$. Let ${\sf P}(\mathcal{C},2\sqrt{N}\sigma)$ denote the packing number, i.e. the maximal number of points $\mathcal{C}$ separated by distance at least $2\sqrt{N}\sigma$; then $\mathsf{N}(\mathcal{C},2\sqrt{N}\sigma)\leq{\sf P}(\mathcal{C},2\sqrt{N}\sigma)$ (see e.g. [42]), and standard volume estimates yields ${\sf P}(\mathcal{C},2\sqrt{N}\sigma)\leq\frac{1}{\min_{t^{N}\in\mathcal{C}}\mu^{\otimes N}(B(t^{N},\sqrt{N}\sigma))}$, proving (51). Moreover, by a volume argument we also have $\mathsf{N}(\mathcal{C},2\sqrt{N}\sigma)\cdot\max_{t^{N}\in\mathcal{C}}\mu^{\otimes N}(B(t^{N},2\sqrt{N}\sigma))\geq\mu^{\otimes N}(\mathcal{C})$. By Lemma 1, we have $\lim_{N\to\infty}\frac{1}{N}\ln\mu^{\otimes N}(\mathcal{C})=0$. Then using (57) with $\sigma$ replaced by $2\sigma$, we obtain (52). ∎
@@
When $|T|<\infty$, it suffices to consider the case of rational $\mu$, since the general case would then follow by a limiting argument. Applying Dudley’s integral (6) to $\mathcal{C}:=\mathcal{C}_{\mu}$ yields
$\mathbb{E}[\max_{z^{N}\in\mathcal{C}}\left<Y^{N},z^{N}\right>]\leq K\int_{0}^{\Delta_{\rm m}}\sqrt{\ln\mathsf{N}(\mathcal{C},2\sqrt{N}\sigma)}2\sqrt{N}\,d\sigma$ (58)
where $\Delta_{\rm m}:=\sup_{N\in\mathcal{I}}\frac{{\rm diam}(\mathcal{C})}{2\sqrt{N}}<\infty$. Now we multiply $1/N$ to (58), take $\mathcal{I}\ni N\to\infty$, and apply Lemma 6 and Lemma 5. We can apply the dominated convergence theorem, since $\frac{1}{N}\sqrt{\ln\mathsf{N}(\mathcal{C},2\sqrt{N}\sigma)}2\sqrt{N}\leq\frac{2}{\sqrt{N}}\sqrt{\ln(n^{N})}=2\sqrt{\ln n}$, which is independent of $N$ and absolutely integrable on $\sigma\in[0,\Delta_{\rm m}]$. Then Theorem 1 for finite $T$ follows from Lemma 6, Lemma 5, and (26). ∎
@@
As in the proof of Theorem 1, we can identify $T=\{e_{1},\dots,e_{n}\}$ as the standard basis vectors, so that $X_{z}=\left<Y,z\right$>. Furthermore, we assume that $|\mathcal{Y}|<\infty$ and $P_{YZ}$ is rational; the general case can then be proved by a limiting argument, since both sides of (2) are continuous with respect to the distribution when $|T|<\infty$. Define
$\mathcal{I}:=\{N\in\mathbb{Z}\colon\textrm{$N>0$, $NP_{YZ}$ is integer}\}.$ (59)
For each $N\in\mathcal{I}$, let $\mathcal{C}_{\mu}$ denote the type class of $\mu$, consisting of all sequences of type $\mu=P_{Z}$. Select
@@
sequences from $\mathcal{C}_{\mu}$ uniformly at random with replacement, and call this random set $\mathcal{A}$. Let $Y^{N}\sim P_{Y}^{\otimes N}$ (independent of this random set). Then we will show that
$\mathbb{E}[\left<Y,Z\right>]\leq\liminf_{\mathcal{I}\ni N\to\infty}\mathbb{E}[\max_{z^{N}\in\mathcal{A}}\frac{1}{N}\left<Y^{N},z^{N}\right$>]. (61)
Indeed, set
@@
and for any $y^{N}\in\mathcal{B}$, define $\mathcal{D}(y^{N}):=\{z^{N}\colon(y^{N},z^{N})\in\mathcal{D}\}$. Standard results about type estimates (immediate from Lemma 2 and Lemma 1) state that:
$\left|\frac{1}{N}\ln|\mathcal{D}(y^{N})|-H(Z|Y)\right|$ $\leq C_{1}\frac{\ln N}{N},\quad\forall y^{N}\in\mathcal{B},$ (64)
$\left|\frac{1}{N}\ln|\mathcal{C}_{\mu}|-H(Z)\right|$ $\leq C_{1}\frac{\ln N}{N},$ (65)
where $C_{1},C_{2}\dots$ in this proof are positive constants independent of $N$. Let $\pi$ be the projection from $\mathbb{R}^{nN}$ to $\mathcal{B}$ (i.e., map to the closest point in $\mathcal{B}$, with ties broken arbitrarily). Define $\mathcal{E}$ as the event that $\mathcal{D}(\pi(Y^{N}))\cap\mathcal{A}=\emptyset$. Then, (64) implies that
$\mathbb{P}[\mathcal{E}]$ $=\left(1-|\mathcal{D}(y^{N})|/|\mathcal{C}_{\mu}|\right)^{L}$ (66)
$\leq\left(1-\exp(-NI(Y;Z)-2C_{1}\ln N)\right)^{L}$ (67)
$\leq\exp(-C_{2}N^{1/3}),$ (68)
where $y^{N}$ is any sequence in $\mathcal{B}$. From the definitions it is clear that
@@
for all $N\in\mathcal{I}$ almost surely. By our assumption of $|T|,|\mathcal{Y}|<\infty$, we have $\left|\max_{z^{N}\in\mathcal{A}}\frac{1}{N}\left<\pi(Y^{N}),z^{N}\right>\right|\leq\max_{y\in\mathcal{Y},z\in T}|\left<y,z\right>|\leq C_{3}$, therefore $\lim_{\mathcal{I}\ni N\to\infty}\mathbb{P}[\mathcal{E}]=0$ and the Cauchy-Schwarz inequality shows that
$\limsup_{\mathcal{I}\ni N\to\infty}\mathbb{E}\left[\left|\max_{z^{N}\in\mathcal{A}}\frac{1}{N}\left<\pi(Y^{N}),z^{N}\right>\right|\cdot 1_{\mathcal{E}}\right]=0.$ (70)
Then (69) implies
$\liminf_{\mathcal{I}\ni N\to\infty}\left[\max_{z^{N}\in\mathcal{A}}\frac{1}{N}\left<\pi(Y^{N}),z^{N}\right>\right]\geq\mathbb{E}[\left<Y,Z\right$>]. (71)
Furthermore,
@@
$\leq C_{4}W_{2}(P_{Y},\widehat{P}_{Y^{N}})\cdot$ (74)
which converges to 0 in expectation [8]. Thus (61) is verified by collecting the bounds above.
@@
$\limsup_{\mathcal{I}\ni N\to\infty}\frac{1}{N}\ln\mathsf{N}(\mathcal{A},2\sqrt{N}\sigma)\leq i_{\mu}(\sigma)\wedge I(Y;Z).$ (75)